### v3

# Strada-LLM: Graph LLM for traffic prediction

Abstract.
Traffic forecasting is pivotal for intelligent transportation systems, where accurate and interpretable predictions can significantly enhance operational efficiency and safety. A key challenge stems from the heterogeneity of traffic conditions across diverse locations, leading to highly varied traffic data distributions. Large language models (LLMs) show exceptional promise for few-shot learning in such dynamic and data-sparse scenarios. However, existing LLM-based solutions often rely on prompt-tuning, which can struggle to fully capture complex graph relationships and spatiotemporal dependencies—thereby limiting adaptability and interpretability in real-world traffic networks.
We address these gaps by introducing Strada-LLM, a novel multivariate probabilistic forecasting LLM that explicitly models both temporal and spatial traffic patterns. By incorporating proximal traffic information as covariates, Strada-LLM more effectively captures local variations and outperforms prompt-based existing LLMs. To further enhance adaptability, we propose a lightweight distribution-derived strategy for domain adaptation, enabling parameter-efficient model updates when encountering new data distributions or altered network topologies—even under few-shot constraints.
Empirical evaluations on spatio-temporal transportation datasets demonstrate that Strada-LLM consistently surpasses state-of-the-art LLM-driven and traditional GNN-based predictors. Specifically, it improves long-term forecasting by 17% in RMSE error and 16% more efficiency. Moreover, it maintains robust performance across different LLM backbones with minimal degradation, making it a versatile and powerful solution for real-world traffic prediction tasks.
spatio-temporal transformers, graph neural networks, traffic prediction, LLM
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: August 03–07, 2025; Toronto, Canada; $c$s: Computing methodologiesccs: Machine learning approaches
1. Introduction
With the development of Intelligent Transportation Systems, spatio-temporal traffic prediction has received increasing attention. It is a key component of advanced urban management systems and is crucial in urban planning, mobility management, and resource allocation. Spatio-temporal prediction involves analyzing urban conditions on various dimensions, including flow, speed, and density, mining their patterns, and predicting trends  (Li et al., 2018). This capability provides a scientific foundation for any urban management department to anticipate and mitigate congestion, implement preemptive restrictions, and enable urban residents to select safer and more efficient travel routes.
Figure 1. An example of the spatio-temporal dependency between the traffic flows for different traffic states ①,②,③. Nodes (a), (b), and (c) are placed in a crowded area of Brussels. The corresponding traffic patterns between 8:$0$ am and 8:$0$ pm are illustrated for nodes (a), (b), and (c). The short-term traffic flow in (c) is altered by the significant impact of neighboring roads, resulting in varying flows.
However, spatio-temporal prediction is a challenging task due to complex spatial and temporal dependencies:
1) Spatial Dependency: The variation in urban patterns is influenced by the topological structure of the urban network. In particular, conditions at upstream locations impact downstream locations through transfer effects, which refers to the impact that a change in one section (e.g., a closure of an area, a change in service timing, or new infrastructure) has on other parts of the network. On the other hand, downstream locations affect upstream locations through feedback effects, which occur when the changes in patterns resulting from transfer effects influence the original cause of the change, creating a loop of cause and effect.
2) Temporal Dependency: Urban conditions change over time, exhibiting periodicity and trends, which are often influenced by factors such as holidays, working hours, and other social events. As illustrated in Figure  1, the strong influence among adjacent locations caused by the flow alongside the edges changes the short-term states. Indeed, location (a) has a steady count trend, and location (b) has short spikes in volume, so the spatio-temporal correlation yielded a different state in location (c) after a couple of hours. Furthermore, urban data sources exhibit a variety of data distributions, posing a significant challenge for finding generalizable models.
In recent years, the emergence of Large Language Models (LLM) has revolutionized various fields due to the generalization capabilities of extracting efficient features from different modalities  (Yuan et al., 2024). Research has been conducted in multiple domains, such as computer vision and natural language processing. Specifically, in  (Yuan et al., 2024), a unified representation space is proposed, namely shared vocabulary, across varied data sources such that the model can learn domain-invariant features. This characteristic is particularly appealing for urban prediction, where data distributions can vary significantly across different regions and times.
Previous works  (Li et al., 2024a, b) have tried to address this problem by casting the graph (and/or the corresponding time-series) to a prompt input. However, this branch of methods, in terms of predictability, is suboptimal because traditional time series forecasting models use numerical inputs and outputs, whereas prompt-based methods transform these into text prompts, which may not be as effective for capturing the characteristics and trends of time series  (Rasul et al., 2024). It leads to a performance gap compared to classical Graph Neural Networks  (Shao et al., 2022). Moreover, the complexity and scalability of the prompt-based approaches are unexplored. For instance, in Spatio-Temporal Graphs (STG) with 1 million nodes, feeding the nodes to an LLM with limited context length is a challenging operation. Consequently, the prompting operation loses the global topological information.
Effective LLM adaptation approach have a noticeable impact on their Zero-Shot performance. There are three prominent approaches to LLM adaptation(Alghisi et al., 2024). Namely, Prompt-Tuning(Qian et al., 2023), Adapter-Layer(Alghisi et al., 2024), Parameter-Efficient Fine-Tuning (PEFT)(Hu et al., 2021). The adapter-layer approach (Alghisi et al., 2024) tends to be more applicable in the multi-task setup for foundational models(Alghisi et al., 2024). In contrast, the effectiveness of Prompt-based approaches and PEFT remains a point of contention in the literature. Both methods have emerged as favored for integrating domain-specific expertise into Large Language Models (LLMs). However, in the realms of existing urban LLMs, UrbanGPT(Li et al., 2024a) and UniST(Yuan et al., 2024) have focused on Prompt-based approaches. Specifically, they include a learnable prompt network (PN) into the LLM. Essentially, the PN is fine-tuned in the adaptation stage. Prompt networks could impose extra complexity and neglect scalable features in the spatio-temporal target domain. Specifically, PN relies on the memory network(Sukhbaatar et al., 2015), which has been proven in graph-based spatio-temporal data to cause exponential error propagation(Barbero et al., 2024). On the other hand, PEFT(Hu et al., 2021) involves modifying a limited set of model parameters—either by targeting existing weights or introducing additional ones—while preserving the bulk of the model’s architecture. PEFT enables the retention of valuable pre-trained feature representations while concurrently tailoring the model to particular domains through targeted adjustments. However, the competence and scalability of PEFT in the LLM-based urban traffic forecasting domain remain unexplored.
To address these issues, we propose a new forecasting method called Strada-LLM, which is a graph-aware LLM for urban traffic prediction. To the best of our knowledge, our model is the first to propose a specialized parameter-efficient LLM model in traffic forecasting. Our contributions are five-fold:
  1. (1)
    Strada-LLM is a probabilistic LLM compliant with classical LLMs and specializes in spatio-temporal prediction. Specifically, Strada-LLM has the capability to predict urban traffic patterns, even when faced with a data distribution that differs from its training data, particularly in the few-shot setting.
  2. (2)
    To the best of our knowledge, Strada-LLM is the first non-prompt-based probabilistic LLM that has lightweight adaptation capability, effectively capturing both local and global spatial dependencies.
  3. (3)
    Strada-LLM is graph-aware, taking spatial dependencies into account by encoding the network graph implicitly and without any prompting. This allows the model to maintain crucial topological information throughout the prediction process.
  4. (4)
    Strada-LLM adopts a lightweight approach to perform domain adaptation. Specifically, adopting low-ranked approaches showcases its effectiveness in adapting to new datasets while maintaining computational efficiency.
  5. (5)
    We provide a comprehensive evaluation and demonstrate superior accuracy compared to existing urban LLMs. We evaluate our approach using numerous real-world datasets. Our results show a reduction in prediction error, ranging from approximately 5% to 18%, compared to baseline methods. This demonstrates the superiority of our graph-aware LLM model in urban forecasting.
The remainder of this paper is organized as follows: Section 2 presents the preliminaries, Section 3 describes the methodology, Section 4 discusses the results, and Section 5 concludes.
2. Preliminaries
Spatio-Temporal Forecasting. In this paper, we aim to predict traffic conditions at a specific timestamp $T{{}^{\prime}}$ based on historical traffic data over period $T$. Particularly, the traffic metric broadly refers to traffic speed, flow, or density. Before further elaborating on our method, we define the following notations. In Strada-LLM, a road network is represented as a graph $G$, consisting of a set of nodes $V$ and a set of edges $E$. Each node presents a traffic measuring sensor, and an edge connects two nodes if they are geographically adjacent. That is, $G=(V,E,A)$ where $V$ is a set of $N$ nodes, $E$ is a set of edges and $A\in\mathbb{R}^{N\times N}$ is the corresponding adjacency matrix. We assume that the topology of the traffic graph $G$ is static. At each timestamp $t$, the graph $G$ is associated with a dynamic feature matrix $\boldsymbol{X}_{t}\in\mathbb{R}^{N\times F}$, where $F$ is the number of traffic metrics.
In this paper, we formulate traffic forecasting as a discrete multi-variate point prediction problem. In particular, Strada-LLM takes the $T$ past observations $\boldsymbol{X}=\left[\boldsymbol{X}_{t_{1}},\boldsymbol{X}_{t_{2}},...,\boldsymbol{X}_{t_{T}}\right]\in\mathbb{R}^{T\times N\times F}$ from $G$ as input and predicts the traffic for horizon $T^{\prime}$, where:
$\boldsymbol{\hat{X}}=\left[\boldsymbol{\hat{X}}_{t_{T+1}},\boldsymbol{\hat{X}}_{t_{T+2}},...,\boldsymbol{\hat{X}}_{t_{T+T^{\prime}}}\right]$}.
3. Methodology
Strada-LLM consists of a Hierarchical Feature Extractor (HFE), an LLM-based backbone, and a T-student distribution head, as shown in Figure 2. The distribution head includes three learnable parameters related to the output: degrees of freedom, mean, and scale, to ensure the relevant parameters remain positive. In the HFE module, a sub-graph extractor and a global graph feature extractor hierarchically learn features while the LLM-based block maps the graph features into the latent spatio-temporal feature space to be learned by the distribution head. In the following sections, we will elaborate on each component.
3.1. Hierarchical Feature Extractor
3.1.1. $k$-hop Sub-Graph Extractor:
To enhance the LLM’s understanding of graph structures while accommodating its limited context length, we tokenize the spatio-temporal traffic signal by taking into account the existing subgraphs of the road network. Overall, the corresponding traffic signals for neighboring nodes will be aggregated together to be tokenized in the further steps. This block is responsible for extracting $k$-hop subgraphs from the input graph. For a node $v\in G$, the k-hop operator (Nikolentzos et al., 2019) is defined as:
(1) $\mathcal{N}_{k}(v)=\left\{u\in V|d(u,v)\leq k\right\},$
where $d(\cdot,\cdot)$ is the hop-based distance function between two nodes. By doing so, each node is equipped with a sub-graph of $G$, including local topological information. As a result, for each node $v$, we concatenate the features of $\mathcal{N}_{k}(v)$, leading to an $N\times T\times(M\times F)$ feature map, where $M$ denotes $|\mathcal{N}_{k}(v)|$, cardinality of neighbors. Similarly, Subgraph-1-WL (Zhou et al., 2024) and MixHop (Abu-El-Haija et al., 2019) can also be used as the k-hop operator, but it is beyond the scope of this research.
3.1.2. Global Graph Embedding
While prompt-based networks are widely used in the literature (Liang et al., 2024), they lose the global structure information (Li et al., 2024b). As demonstrated by  (Trillos et al., 2021), in order to keep the global graph structure information, we compute Laplacian embeddings. The Laplacian operation leverages the eigenvectors of the complete graph and creates unique positional encodings (Vaswani et al., 2023) for each node. First, we define the graph Laplacian matrix  $\textbf{L}=D-A$ where $D=\text{diag}(d_{1},\ldots,d_{N})$ is the degree matrix. We then compute the eigendecomposition of the normalized Laplacian as follows $L_{\text{norm}}=D^{-1/2}\textbf{L}D^{-1/2}=\textbf{U}\Lambda\textbf{U}^{T}$. Then we compute global embedding by non-linear mapping of U matrix. Reader can refer to Section 4.4.3 to discern the contribution of this representation extraction on the LLM forecasting ability. We denote $F^{\prime}$ as the embedding dimension.
3.1.3. Lag Extractor
Mathematically, this lagging operation can be expressed as a mapping $\boldsymbol{x}_{t}\longmapsto\boldsymbol{\lambda_{t}}\in\mathbb{R}^{|\mathcal{H}|\times F}$, where $\mathcal{H}=\left\{h_{1},h_{2},...,h_{H}\right\}$ represents the set of historical lag indices. In this formulation, $\lambda_{t}[j]=x_{t-\mathcal{H}[j]}$, meaning that $\lambda_{t}[j]$ corresponds to the value of $x$ at $h_{j}$ time steps before $t$, as specified by the $j$-th element of $\mathcal{H}$.
In our approach, we extend beyond simple lag feature extraction by employing a structured tokenization scheme to convert spatio-temporal slices of the road network graph into discrete tokens for the LLM-based backbone. Specifically, for each time step $t$, we first derive both global and local (k-hop) graph embeddings. These representations are then segmented within a sliding window, yielding embeddings that encapsulate the node-level or sub-graph-level features alongside their temporal contexts. Crucially, we add positional embeddings to each token so that the masked Transformer decoder can effectively attend to different time steps in a sequence-like manner. While lag features help capture short-term correlations, the tokenization scheme formalizes how each time window becomes a coherent input unit, thus ensuring the LLM’s attention mechanisms can leverage meaningful traffic patterns and dependencies across time and space.
Therefore, to generate lag features for a specific context-length window $x_{1:C}$, where $C$ is the length of the context window, it is necessary to consider an expanded window that includes $H$ additional historical data points. Besides the lagged features, we incorporate date-time features covering all temporal features in our dataset, including second-of-minute, hour-of-day, and so forth, up to quarter-of-year, based on the time index $t$. It is important to note that the main purpose of these date-time features is to enrich the information set. In any time series, all but one of the date-time features will stay the same from one-time step to another, allowing the model to intuitively understand the time series frequency. In the end, the HFE block’s output is a tensor in shape $\mathbb{R}^{C\times(1+M)\times|\mathcal{H}|\times(F+F^{\prime})}$.
3.2. LLM-based-Backbone
Our backbone architecture is inspired by Mistral (Jiang et al., 2023), a decoder-only transformer architecture. The extracted tokens by the sub-graph extractor are transformed by a shared linear projection layer that maps the features to the hidden dimension of the attention module. Inspired by (Touvron et al., 2023), Strada-LLM includes pre-normalization by utilizing the RMSNorm (Zhou et al., 2021) and Rotary Positional Encoding (RoPE) (Su et al., 2023) for the representation of query and key blocks at each attention layer, similar to the approach described for Mistral (Jiang et al., 2023).
After passing through the causally masked transformer layers, the proposed method incorporates Flash-Attention-2 (Dao, 2023) to predict the parameters of the forecast distribution for the next timestamp, denoted as $\phi$. These parameters are generated by a parametric distribution head, as shown in Figure 2. The objective is to minimize the negative log-likelihood of the predicted distribution across all prediction times.
At the inference stage, given a time series with the size of at least $H+C$, a feature vector can be formed and supplied to the model to determine the distribution of the subsequent timestamps. In this fashion, we can obtain many simulated future trajectories up to our chosen prediction horizon $T^{\prime}$ via auto-regressive decoding (Bernardo et al., 2007). In doing so, uncertainty intervals and confidence scores could also be valuable for downstream decision-making tasks.
Figure 2. The proposed Strada-LLM architecture. The core novelty lies in the K-hop sub-graph extractor module, which takes lag features in both the spatial and temporal dimensions.
3.3. Domain Adaptation
In this section, we describe the techniques we adopt to adapt the Strada-LLM model to a new dataset, which might contain a graph of a new city. One of the important aspects of such a model is that it should be domain-agnostic. In the rapidly evolving landscape of deep learning, domain adaptation has emerged as a crucial technique for enhancing the generalizability of models across different but related domains. More specifically, domain adaptation focuses on the ability of a model trained in one domain (source) to adapt and perform well in another domain (target) with potentially different distributions. A prominent approach within this realm is low-rank adaptation (Hu et al., 2021), which aims to refine pre-trained models by adapting their parameters in a computationally efficient manner; indeed, a nearly linear operation will replace the $\mathcal{O}(n^{2})$ operation. This technique is particularly beneficial when working with LLMs, which, despite their impressive performance and versatility, pose significant challenges regarding computational resources and training time. In this paper, we adapt to the new distribution by tuning the distribution head, query, key, and value attention blocks.
(a) Strada-LLM (our method)
(b) UniST(Yuan et al., 2024)
Figure 3. Conceptual LLM adaptation approaches comparison
3.3.1. Low-Rank Matrix Adaptation
By leveraging domain adaptation strategies and Low-Rank Matrix Adaptation (LoRA) for fine-tuning, researchers and practitioners can effectively bridge the gap between domains, ensuring that powerful deep learning models maintain their efficacy across diverse applications (Filatov and Kindulov, 2023). One of the purposes of this research is to propose a LoRA-based method for the temporal graph forecasting problem. As illustrated in Figure 3, we apply LoRA for the LLM fine-tuning process such as spatio-temporal patterns of the target data being learned. The query, key, and value matrices will be fine-tuned in a low-rank approximation manner. Specifically, according to Figure 3, suppose the query, key, and value tensors are computed as follows:
(2) $\small q=W_{q}h,k=W_{k}h,v=W_{v}h,$
where $h$ is the output of the last layer of the LLM backbone with dimension $\mathbb{R}^{d\times k}$. $W_{q},W_{k}$ and $W_{v}$ are the query, key, and the projection weight matrices respectively. The low-rank matrices for fine-tuning the forward step can be defined as:
(3) $\small\begin{split}\small q=W_{q}h+\Delta W_{q}h=W_{q}h+B_{q}A_{q}h\\ k=W_{k}h+\Delta W_{k}h=W_{k}h+B_{k}A_{k}h\\ v=W_{v}h+\Delta W_{v}h=W_{v}h+B_{v}A_{v}h\end{split}$
where $h$ is the output of the previous layer, the low-ranked approximated update matrices $B_{q},B_{k},B_{v}$ are of dimensions $\mathbb{R}^{d\times r}$, $A_{q},A_{k},A_{v}$ are of dimensions $\mathbb{R}^{r\times k}$, and where the rank hyperparameter $r\ll min(d,k)$. By introducing these matrices, online inference latency reduction is achieved (Hu et al., 2021). For the initialization of the matrices $A$ we choose Gaussian initialization while the $B$ matrices are initialized to zero (Hu et al., 2021). Following this approach enables us to perform domain adaptation for different datasets consisting of diverse distributions. As shown in Figure 3(a), our adaptation approach is conceptually different than prompt-based models like UniST (Yuan et al., 2024). Based on the distribution-driven alignment loss function in the adaptation phase, global graph embedding, query, key, value projections, and distribution head are targeted to be tuned in the low-rank fashion. As reader can see in the experiment section 4.4.4, tuning the mentioned blocks is more lightweight than tuning prompt-based models like Figure 3(b). One might critique not applying low-rank tuning for other modules like Global Graph Embedding in Figure 3(a), which is addressed in Appendix  A.4.
Recent theoretical work shows that domain alignment can be bounded using the optimal transport distance functions (Yao et al., 2020). For instance, unsupervised distance adaptation approaches (Ramesh et al., 2024) have been adopted in modern LLMs. As shown in Figure  3(a), In Strada-LLM, the distribution head will be tuned. So, as another novel pillar of our model, during the adaptation phase, we minimize Kullback–Leibler divergence $D_{KL}(.)$ distance between multivariate distribution heads. As a matter of tractability, given a bounded $k$-variate $t$-distribution, the distance function can be derived as:
$D_{KL}(f_{d}(\cdot|\mu,\Sigma,\nu)||f_{d}(\cdot|\mu,\Sigma,\nu^{\prime}))=$
$=D_{KL}(f_{d}(\cdot|0,\mathbf{I},\nu)||f_{d}(\cdot|0,\mathbf{I},\nu^{\prime}))$
$=\int_{\mathbb{R}^{n}}f_{d}(\mathbf{x}|0,\mathbf{I},\nu)\log\frac{f_{d}(\mathbf{x}|0,\mathbf{I},\nu)}{f_{d}(\mathbf{x}|0,\mathbf{I},\nu^{\prime})}d\mathbf{x}$
Inspired by  (Villa and Rubio, 2018), this Kullback Leibler divergence can be written as:
(4) $D_{KL}(f_{d}(\cdot|\mu,\Sigma,\nu)||f_{d}(\cdot|\mu,\Sigma,\nu^{\prime}))=$
$\log\frac{K(d,\nu)}{K(d,\nu^{\prime})}-\frac{\nu+d}{2}\left[\Psi\left(\frac{\nu+d}{2}\right)-\Psi\left(\frac{\nu}{2}\right)\right]+$
$\frac{\nu^{\prime}+d}{2}K(d,\nu)\frac{\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2})}\int_{0}^{\infty}\left(1+\frac{t}{\nu}\right)^{-\frac{\nu+d}{2}}t^{\frac{d}{2}-1}\log\left(1+\frac{t}{\nu^{\prime}}\right)dt$
$\coloneqq{\mathcal{L}_{alignment}}$
where, $\small{K(d,\nu)=\frac{\Gamma(\frac{\nu+d}{2})}{\Gamma(\frac{\nu}{2})\sqrt{(\pi\nu)^{d}}}}$. The proof unfolds in Appendix A.2.
3.4. Loss Function
The Negative Log-Likelihood (NLL) is a loss function commonly used in probabilistic time-series forecasting. It is calculated as the negative of the log of the probability of the true value given the predicted value. The formula for negative log-likelihood is:
(5) $\small{\text{NLL}=-\sum_{x\in\mathcal{D}}log(P(y|x))},\mathcal{L}_{total}=\lambda_{1}NLL+(1-\lambda_{1})\mathcal{L}_{alignment}$
where $y$ is the true value, $x$ is the input data, $\mathcal{D}$ is the training set, $P(y|x)$ is the predicted conditional probability of $y$ given $x$, and $\lambda_{1}$ is the loss component hyperparameter. In multivariate modeling, $NLL$ can be modeled differently. In Strada-LLM, we estimate the joint distribution, but further experiments are discussed in the ablation study 4.4.5.
4. EXPERIMENTS
Figure 4. Comparison of unified benchmark for different LLMs.
4.1. Datasets Details
In this section, we assess the Strada-LLM model’s predictive capabilities using eight real-world datasets. The description of these datasets is presented in Table 1. Most of these datasets pertain to traffic speed. For the sake of simplicity, traffic speed is utilized as the primary traffic metric in our experimental analysis; thus one can presume $F=1$. Detailed training strategies, e.g., hyperparameters and experimental setup, are explained in Appendix A.5.
Table 1. Dataset Description
Dataset Data Type Number of Nodes Region Time Steps Pretraining Datasets PeMS03 (Fang et al., 2021) Traffic Speed 170 California 26208 PeMS04 (Fang et al., 2021) Traffic Speed 307 California 16992 PeMS07 (Fang et al., 2021) Traffic Speed 883 California 26208 PeMS08 (Fang et al., 2021) Traffic Flow 170 California 17856 Few-Shot METR-LA (Fang et al., 2021) Traffic Speed 207 California 34272 PEMS-Bay (Fang et al., 2021) Traffic Speed 325 California 52116 PeMS07(M) (Fang et al., 2021) Traffic Speed 228 California 12672 Brussels (Moghadas et al., 2024) Traffic Count 207 Brussels 12672 Zero-Shot Crowd (Yuan et al., 2024) Pedestrian Count $16\times 20$ Nanjing 9420 PeMS07(M) (Fang et al., 2021) Traffic Count 228 California 12672 Brussels (Moghadas et al., 2024) Traffic Count 207 Brussels 12672
4.2. Baselines
We compare the performance of the Strada-LLM model with spatio-temporal methods, GMAN  (Zheng et al., 2020), MTGNN  (Wu et al., 2020), GTS  (Shang et al., 2021), STEP  (Shao et al., 2022), and LLM-based methods which either are point-wise forecasters such as FlashST (Li et al., 2024b), UniST (Yuan et al., 2024) or probabilistic forecasters such as Lag-llama(Rasul et al., 2024). The details of the baseline methods are explained in Appendix A.1. To compare with the baselines, we adopt the MAE, RMSE, and MAPE prediction error metrics (Li et al., 2024b). Due to Strada-LLM being a probabilistic model, we also compare with Lag-Llama (Rasul et al., 2024) using the Continuous Ranked Probability Score (CRPS) metric, defined as:
$\small{CRPS(F,y)=\int_{-\infty}^{\infty}(F(x)-H(x-y))^{2}dx,}$
where $F(x)$ is the cumulative distribution function (CDF) of the forecasted distribution, $y$ is the observed value, and H(xy)H(x-y) is the Heaviside step function, which is 0 when $x<y$ and 1 when $x\geq y$.
4.3. LLM Forecasting Results
4.3.1. LLM Short-Term Forecasting
The short-term traffic prediction accuracy was evaluated across four datasets: PeMS-Bay, METR-LA, PEMS07(M), and Brussels. The last dataset is private. Strada-LLM, as an LLM should be benchmarked jointly for all datasets.
The prediction results for the datasets METR-LA and PEMS-Bay are reported in Tables 23, respectively. The ability to perform forecasting and being robust to new domains are important qualities for an LLM. As shown in Tables 23, the proposed LLM has competitive performance compared to STEP (Shao et al., 2022), which is the corresponding transformer-based fully-supervised method. Readers can refer to Appendix 10 about robustness analysis. On the other hand, compared to the best prompt-tuning method FlashST (Li et al., 2024b), shows that taking the global structure of the graph into account brings forth a definite advantage. Moreover, although Strada-LLM and STEP (Shao et al., 2022) are both transformer-based models, the introduction of normalization blocks like RMSNorm (Zhou et al., 2021) and Rotary Positional Encoding (RoPE) differentiates them. Additionally, while STEP (Shao et al., 2022) is an encoder-decoder model, Strada-LLM is a decoder-only model, which makes it optimized in terms of performance for inference (Fu et al., 2023). As we can see in Tables 23, the proposed Strada-LLM manages to outperform the baseline models on the aforementioned datasets in the longer prediction horizon more significantly. Specifically, on the PeMS-Bay dataset, our proposed model achieves an RMSE of 3.94 for 1-hour prediction, corresponding to an improvement of approximately $16\%$ compared to the prompt-tuning baseline (Li et al., 2024b).
Table 2. Traffic short-term forecasting comparison with the state-of-the-art methods on the METR-LA dataset
Baselines 15 min 30 min 60 min
[…]
Supervised
GMAN (Zheng et al., 2020) 2.80 5.55 7.41 3.12 6.49 8.73 3.44 7.35 10.07
MTGNN (Wu et al., 2020) 2.69 5.18 6.88 3.05 6.17 8.19 3.49 7.23 9.87
GTS (Shang et al., 2021) 2.67 5.27 7.21 3.04 6.25 8.41 3.46 7.31 9.98
STEP(Shao et al., 2022) 2.61 4.98 6.6 2.99 5.97 7.96 3.37 6.99 9.61
LLM
FlashST (Li et al., 2024b) 2.84 5.57 7.45 3.19 6.43 9.04 3.60 7.44 10.68
UniST (Yuan et al., 2024) 2.89 5.63 7.49 3.26 6.51 9.09 3.65 7.52 10.79
Strada-LLM(ours) 2.45 4.71 6.42 2.97 5.84 7.87 3.35 6.73 9.47
Table 3. Traffic short-term predictions comparison with the state-of-the-art methods on PEMS-Bay dataset
Baselines 15 min 30 min 60 min
[…]
Supervised
GMAN (Zheng et al., 2020) 1.34 2.91 2.86 1.63 3.76 3.68 1.86 4.32 4.37
MTGNN (Wu et al., 2020) 1.32 2.79 2.77 1.65 3.74 3.69 1.94 4.49 4.53
GTS (Shang et al., 2021) 1.34 2.83 2.82 1.66 3.78 3.77 1.95 4.43 4.58
STEP(Shao et al., 2022) 1.26 2.73 2.59 1.55 3.58 3.43 1.79 4.20 4.18
LLM
FlashST (Li et al., 2024b) 1.37 2.96 2.88 1.72 3.98 3.89 2.02 4.69 4.79
UniST (Yuan et al., 2024) 1.42 3.01 2.91 1.78 3.99 3.91 2.07 4.72 4.81
Strada-LLM (Ours) 1.28 2.77 2.61 1.53 3.14 3.37 1.76 3.94 4.09
Table 4. Long-Term prediction experiment on the PEMS07 (M) and Brussels dataset ($p\text{-value}<0.05$)
[…]
4.3.2. LLM Long-Term Forecasting
Followed by the setup of  (Yuan et al., 2024), we experiment with long-term forecasting for horizons 90, 120, and 150 minutes. The long-term traffic prediction accuracy was evaluated across two datasets: Crowd and PEMS07(M). We report the performance in Table 4. Consistently, Strada-LLM outperforms the LLM-based baselines. The successful performance can be attributed to the Graphical modeling of spatial covariates. Further visualizations are presented in the Appendix A.3.
Table 5. Few-shot learning experiment on the PEMS07(M) and Brussels dataset($p\text{-value}<0.05$)
[…]
4.3.3. LLM Zero-Shot Forecasting
The zero-shot traffic prediction accuracy was evaluated across three datasets: Crowd, Pems07(M), and Brussels. Table  7 presents the results of zero-shot capability of Strada-LLM compared to the other LLMs. Further visualizations are presented in the Appendix A.3. Strada-LLM is able to capture the spatio-temporal dependencies between connected nodes and prediction correlated trends, which is neglected by UniST (Yuan et al., 2024). To demonstrate the generalizability to the other urban datasets, Strada-LLM’s performance on the Crowd dataset (Yuan et al., 2024) is demonstrated in Figure 5, which outperforms other LLM baselines.
15min30min60min6.56.5777.57.5886.536.536.876.877.177.176.866.867.057.057.317.317.027.027.127.127.637.63HorizonsMAEOursUniST(Yuan et al., 2024)FlashST(Li et al., 2024b)
Figure 5. Zero-shot performance comparison on Crowd dataset.
4.3.4. LLM Probabilistic Forecasting
Table 6 reports the CRPS metric, likewise Lag-llama(Rasul et al., 2024), it has been implemented as the mean quantile loss. We observe that Strada-LLM consistently achieve strong few-shot and few-shot performance, obtaining either the best or second-best results for the Pems07 and Brussels datasets.
Table 6. Probabilistic forecasting experiment on the PEMS07(M) and Brussels dataset
PEMS07(M) Brussels
Baselines $\downarrow CRPS\downarrow$
15 min 30 min 60 min 15 min 30 min 60 min
Lag-llama (Rasul et al., 2024) 3.95 4.13 4.92 5.07 5.77 5.91
Strada-LLM(ours) 3.13 3.67 4.03 4.93 5.10 5.31
Table 7. Zero-shot learning experiment on the PEMS07(M) and Brussels dataset($p\text{-value}<0.05$)
[…]
4.4. Ablation Study
Table 8. Domain adaptation comparison for demonstrating selection of effective source dataset
Source $\rightarrow$ Target $Top^{k}$ Fine-tuning Ours Fully-Finetune Fully-Supervised
$downarrow MAE\downarrow RMSE\downarrow$ $downarrow MAE\downarrow RMSE\downarrow$
Metr-LA Pems-BAY 3.31 5.76 2.64 4.83 2.01 3.89 1.54 3.53
PeMS04 Pems-BAY 2.73 5.02 1.96 4.15 1.67 3.67 1.54 3.53
Table 9. The computation cost on the PEMS07(M) with the batchsize=64.
Method Parameters(M)
Trainable
Parameters(M)
Trainable
[…]
FlashST(Li et al., 2024b) 64.02 1.01 1.58
UniST(Yuan et al., 2024) 30.00 0.71 2.39
Strada-LLM (ours) 40.00 0.52 1.32
4.4.1. Domain Adaptation
We examine the adaptation capabilities of Strada-LLM by evaluating the gap when different datasets are used as source and target domains. Specifically, we trained Strada-LLM on one dataset (e.g., PeMS-04) and adapted to another (e.g., METR-LA) to assess its generalization across different types of traffic patterns. Table 8 presents the results for the domain adaptation process when pre-trained on a source dataset and adapted to the corresponding target dataset. The results presented in Table 8 indicate that there is a performance drop when the model is applied to a different target domain. This table can also address selecting efficient pre-training datasets to gain transferability advance for the targeted LLM. Take Pems-Bay as an example, selecting Pems04 as the pertaining dataset for LLM ends up the 14.07% superior performance in the adaptation phase compared to Metr-LA, because of the inherent closer distribution. Detailed ablated studies about the performance of LoRA in terms of the rank parameter are discussed in Appendix A.6.
Figure 6. Latent space embedding of the last layer (before the distribution head) visualized in a 2D space.
4.4.2. Model Interpretation
To better understand Strada-LLM, we visualize the latent space of the LLM representation for different datasets. By doing so, the disparity of dataset distributions and the reason for the difficulty of adapting to PeMS04 are revealed. The visualization is shown in Figure 6 which is obtained with T-SNE (Cai and Ma, 2022) in 2D dimensions. Our variational posterior distribution $P(\theta\mid X)$ can truly depict the underlying distribution.
4.4.3. Scalability of K-hop structure
Our experimental analysis explores Strada-LLM’s sensitivity to different neighborhood radii by examining performance across multiple k-hop values ($k\in\{1,3,5\}$). The results in Table 10 demonstrate a trade-off between prediction accuracy and neighborhood size, with larger radii yielding lower error rates. Comparative analysis presented in Table 6 shows that Strada-LLM achieves superior performance over Lag-Llama (Rasul et al., 2024), which did not take neighbors into account, primarily due to our model’s incorporation of spatial dependencies through the K-hop block structure, a feature absent in the baseline approach.
Table 10. Feature Scalability Analysis Benchmark
[…]
11 2.96 5.34 7.35
33 2.60 4.86 6.51
55 2.63 4.89 6.54
4.4.4. Computation Cost
We evaluate the performance of our proposed LoRA-based approach compared to existing LLMs. Results are demonstrated in Table 9. As can be seen, the lower error comes with the cost of involving more parameters. However, in contrast to UniST (Yuan et al., 2024), Strada-LLM’s adaptation outperforms the compression rate, from 1.58(%) to 1.32(%), which demonstrates its efficiency. For real-time applications, more lightweight models are being demanded, which remains the future direction. Performance over parameter analysis is illustrated in Appendix A.8.
4.4.5. Distribution modeling
We evaluate our model’s flexibility by deploying it across multiple LLM architectures. So, we report the performance in Table 11. Consistently, Strada-LLM outperforms the LLM-based baselines. The successful performance can be attributed to the Graphical modeling of spatial covariates.
Table 11. Ablation study on distribution modeling for PEMS07(M)
[…]
Univariate 2.81 5.63 6.81
Independent 2.69 5.37 6.73
Joint 2.60 4.86 6.51
4.4.6. LLM Cross-Compatibility
We evaluate our model’s flexibility by deploying it across multiple LLM architectures. Specifically, we utilized Mistral and Llama 2 as base models to assess the adaptability of our proposed method. The results are shown in Table 12.
Table 12. Ablation study on LLM backbone adaptability for PEMS07(M)
[…]
Llama2 (Rasul et al., 2024) 2.83 5.66 6.89
Mistral (Jiang et al., 2023) 2.60 4.86 6.51
5. Conclusion
This research proposes a probabilistic-based LLM for traffic forecasting called Strada-LLM, which injects the local graph structure implicitly into the input tokens. We utilize a subgraph extraction procedure to inject the graph structure into the transformer token inputs. On the one hand, we adopt a probabilistic transformer to predict the traffic data by sampling from the underlying distribution. On the other hand, we utilize a low-rank method for the transfer learning task. Consequently, we showcase its success in adapting to datasets with significant differences compared to the datasets used for pre-training. Finally, Strada-LLM is compliant to a variety of known LLM backbones. In summary, the Strada-LLM model successfully captures the spatial and temporal features from traffic data so that they can be applied to other spatio-temporal tasks. A potential avenue for future research could involve evaluating the expressiveness of the distribution head.
Appendix A Appendix
A.1. Baselines
We compare the performance of the Strada-LLM model with the following baseline methods:
  • GMAN  (Zheng et al., 2020): The gating fusion mechanism redefines the spatio-temporal attention block.
  • MTGNN  (Wu et al., 2020): Alternating use of graph convolution and temporal convolution modules.
  • GTS  (Shang et al., 2021): The underlying graph structure is learned among multiple time series, and the corresponding time series is simultaneously predicted with DCRNN (Li et al., 2018).
  • STEP  (Shao et al., 2022): Adopts the graph wavenet model (Wu et al., 2019), integrated into the transformer backbone and pre-training scheme.
  • FlashST (Li et al., 2024b), UniST (Yuan et al., 2024): The data distribution shift is learned by leveraging prompt-tuning.
  • Lag-Llama (Rasul et al., 2024): They used known LLM backbones for unified probabilistic time-series forecasting.
To compare with the baselines, we adopt MAE, RMSE, and MAPE prediction error (Li et al., 2024b).
A.2. Proof of Equation LABEL:kl_formula
Let $f_{d}({\bf x}\mid\boldsymbol{\mu},\boldsymbol{\Sigma},\nu)$ be a multivariate $t$, of dimension $d$, with location vector $\boldsymbol{\mu}$, scale matrix $\boldsymbol{\Sigma}$ and $\nu$ degrees of freedom. The aim is to define an objective prior for the parameter $\nu$. For simplicity in the notation, we will write $f_{d,\nu}=f_{d}({\bf x}\mid\boldsymbol{\mu},\boldsymbol{\Sigma},\nu)$, for $\nu=1,\ldots,\nu_{\max-1}$, and $f_{d,\nu_{\max}}=N_{d}({\bf x}\mid\boldsymbol{\mu},\boldsymbol{\Sigma})$, with
$N_{d}({\bf x}\mid\boldsymbol{\mu},\boldsymbol{\Sigma})=\frac{1}{\sqrt{(2\pi)^{d}|\boldsymbol{\Sigma}|}}\exp\left\{-\frac{1}{2}({\bf x}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}^{-1}({\bf x}-\boldsymbol{\mu})\right\},$
where in this case $\boldsymbol{\mu}$ is the vector of means and $\boldsymbol{\Sigma}$ is the covariance matrix. The prior for $\nu$ here discussed depends on the Kullback–Leibler divergence between two multivariate densities. In particular, for $\nu=1,\ldots,\nu_{\max-1}$, the prior is based on the Kullback–Leibler divergence between two multivariate $t$ densities, which differ only in the number of degrees of freedom. The divergence between two $d$-variate $t$ densities, $f_{d,\nu}$ and $f_{d,\nu^{\prime}}$, is given by
$D_{KL}(f_{d}(\cdot\mid\boldsymbol{\mu},\boldsymbol{\Sigma},\nu)\mid\mid f_{d,}(\cdot\mid\boldsymbol{\mu},\boldsymbol{\Sigma},\nu^{\prime}))$
$=$ $D_{KL}(f_{d}(\cdot\mid{\bf 0},{\bf I},\nu)\mid\mid f_{d}(\cdot\mid{\bf 0},{\bf I},\nu^{\prime}))$
$=$ $\int_{{\mathbb{R}}^{n}}f_{d}({\bf x}\mid{\bf 0},{\bf I},\nu)\log\dfrac{f_{d}({\bf x}\mid{\bf 0},{\bf I},\nu)}{f_{d}({\bf x}\mid{\bf 0},{\bf I},\nu^{\prime})}\,d{\bf x}$
$=$ $\int_{{\mathbb{R}}^{n}}K(d,\nu)\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu}\right)^{-\frac{\nu+d}{2}}\log\dfrac{K(d,\nu)\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu}\right)^{-\frac{\nu+d}{2}}}{K(d,\nu^{\prime})\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu^{\prime}}\right)^{-\frac{\nu^{\prime}+d}{2}}}\,d{\bf x}$
(8) $=$ $\log\dfrac{K(d,\nu)}{K(d,\nu^{\prime})}-\dfrac{\nu+d}{2}{\mathbb{E}}_{d,\nu}\left[\log\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu}\right)\right]+$
$\dfrac{\nu^{\prime}+d}{2}{\mathbb{E}}_{d,\nu}\left[\log\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu^{\prime}}\right)\right],$
where
$K(d,\nu)=\dfrac{\Gamma\left(\dfrac{\nu+d}{2}\right)}{\Gamma\left(\dfrac{\nu}{2}\right)\sqrt{(\pi\nu)^{d}}},$
and ${\mathbb{E}}_{d,\nu}$ represents the expected value with respect to $f_{d}(\cdot\mid{\bf 0},{\bf I},\nu)$. More specifically,
${\mathbb{E}}_{d,\nu}\left[\log\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu}\right)\right]=\Psi\left(\dfrac{\nu+d}{2}\right)-\Psi\left(\dfrac{\nu}{2}\right),$
where $\Psi$ is the digamma function. For the second expectation in (A.2), we use Lemma 2 in (Zografos, 1999) to obtain the following expression after a change of variables in terms of spherical coordinates
$\mathbb{E}_{d,\nu}\left[\log\left(1+\frac{\mathbf{x}^{\top}\mathbf{x}}{\nu^{\prime}}\right)\right]$ $=$ $K(d,\nu)\frac{\pi^{\frac{d}{2}}}{\Gamma\left(\frac{d}{2}\right)}\times$
$\int_{0}^{\infty}\left(1+\frac{t}{\nu}\right)^{-\frac{\nu+d}{2}}t^{\frac{d}{2}-1}\log\left(1+\frac{t}{\nu^{\prime}}\right)dt$
which only requires one-dimensional numerical integration, regardless of the dimension $d$.
Figure 7. The visualization of $k$-hop multivariate zero-shot prediction for Pems07(M) dataset for long-term horizon
(a) Node 6
(b) Node 13
(c) Node 48
Figure 8. Strada-LLM traffic prediction in the capturing volatility scenarios for three (not-connected) different nodes exist in Pems07(M) dataset.
A.3. Visualizations
To better understand and assess our model visually, we present additional illustrations in this section. The visualization of long-term forecasting in zero-shot setting for Pems07(M) dataset is shown in Figure 7. Figure 7 is an instance of multivariate probabilistic forecasting in which the dark blue nodes are part of the k-hop extracted subgraph. A critical aspect of time-series prediction algorithms is their ability to capture volatility importance - the varying degrees of uncertainty and rapid changes in the data over time. By incorporating volatility measures, these algorithms can better understand periods of high traffic turbulence versus relative stability, leading to more nuanced and accurate predictions. Strada-LLM is also able to capture volatility existing in Pems07(M) dataset, which is shown in Figure  8.
A.4. Gradient Analysis
As mentioned in section  3.3, the low-rank adaptation has not been utilised in the Graph Embedding module, which causes model performance degradation. To explain the reason, the norm and variance magnitude of the gradients of two MLP layers inside the module are plotted in Figure 9. As shown in Figure 9(b), according to the heterogeneous gradient backpropagation of these layers, e.g., one MLP’s variance is 10 times the other MLP layer, and the adoption of the same low-rank adaptation causes performance degradation. Proposing an efficient low-rank adaptation in this use case will remain an ongoing research direction.
(a) Comparison of $l_{2}$ gradient norms between the first layer (left) and the second layer (right) across training iterations.
(b) Gradient $\sigma^{2}$ variance comparison between LaplacianPE1 (left) and LaplacianPE2 (right) layers during training.
Figure 9. Training dynamics comparison between two graph Laplacian positional encoding layers: gradient norms and variances over iterations.
A.5. Training Strategy
A.5.1. Augmentation Techniques
We adopted the time series augmentation techniques FreqMixFreq\-Mix(Chen et al., 2023) and FreqMaskFreq\-Mask (Chen et al., 2023) to prevent over-fitting. Our training framework comprises two distinct stages: pre-training and fine-tuning. During pre-training, we leverage a diverse combination of datasets: $D_{base}$ for general language understanding, $D_{domain}$ for domain-specific knowledge, and $D_{task}$ for task-oriented capabilities. These datasets are processed in a curriculum learning manner, starting with $D_{base}$ to establish foundational knowledge, followed by $D_{domain}$ to inject domain expertise, and finally $D_{task}$ to enhance task-specific performance. To maintain balanced exposure across datasets while optimizing computational efficiency, we employ a dynamic batch construction strategy where each training batch $B_{t}$ consists of samples drawn from multiple datasets with proportional representation: $\alpha_{1}$% from $D_{base}$, $\alpha_{2}$% from $D_{domain}$, and $\alpha_{3}$% from $D_{task}$, where $\sum_{i}\alpha_{i}=1$. This balanced approach ensures the model develops robust general capabilities while maintaining strong performance on domain-specific tasks. In the fine-tuning stage, we exclusively focus on $D_{task}$ with smaller learning rates and specialized optimization techniques to prevent catastrophic forgetting of pre-trained knowledge while adapting to target objectives.
A.5.2. Experimental Setup
The hyperparameters of the Strada-LLM model mainly include: learning rate, batch size, training epochs, the number of layers, context length, embedding dimension per head, number of heads, rope scaling, and number of parallel samples for greedy decoding. We adopt Adam as the optimizer (Kingma, 2014). In the experiment, we follow the weight decay 1e81e-8 and set the learning rate to 0.001 (Kingma, 2014), the batch size to 32, and the training epochs to 150 for each dataset. We set $k=3$ in the sub-graph extractor. Furthermore, based on the fact that Strada-LLM is a probabilistic model, we take samples and compute medians to compare with point-wise models. The number of samples is set to $100$. In terms of complexity, our model contains 16 million parameters. The experiments are conducted on 2 Nvidia RTX 4090 GPUs.
Table 13. LoRA rank benchmark results on PeMS04 dataset
[…]
22 24.76 29.37 12.80
44 22.12 27.58 12.09
88 21.87 26.62 11.8
A.6. Rank Analysis
We evaluate the performance of our LoRA adoption with different widths. To do so, we try various values for $r\in(2,4,8)$ and the results are demonstrated in Table 13. As can be seen, the lower error comes with the cost of involving more parameters. For real-time applications, the lower bandwidths are more applicable.
A.7. Perturbation Analysis and Robustness
To evaluate the robustness of the Strada-LLM model against noise, we conduct perturbation analysis experiments. We added two types of common random noise to the validation data during the experiment. Random noise obeys the Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ where $\sigma\in(0.2,0.4,0.8,1,2)$. Then, we normalize the values of the noise matrices to be between 0 and 1. The results are shown in Figure 10, where the horizontal axis represents $\sigma$, the vertical axis represents the error, and different colors indicate different evaluation metrics. This benchmark is evaluated for the dataset PeMS04 but other datasets follow the same pattern. According to this experiment, our model is robust to Gaussian noise.
Figure 10. Evaluation metrics for the perturbation analysis. Different $\sigma$ values were examined to test the model’s robustness.
Figure 11. Strada-LLM scalability analysis compared to existing LLMs.
A.8. Scalability Comparison
Our empirical analysis reveals a nuanced relationship between model size and performance, characterized by diminishing returns beyond certain parameter thresholds. As illustrated in Figure 11, our model performs superior on the similar parameter capacity on the Crowd dataset.